The steam point and the ice point of mercury thermometer are wrongly marked as 97degree centigrade and 2 degree centigrade respectively .the correct temperature read by this thermometer is

Grade:11

1 Answers

Arun
3 years ago
If the options are given we can solve the question
 
X-2/92-2=C/100
 
Here assume x as 20 from options
 
20-2/90=C/100
 
C=20C